# with Mixture of Frequency Adaptation

### v3

@@ 4 Experiments
CLIP [54] (ViT-Large) Frozen 0.0M 43.7 44.4 44.1
Full Fine-Tune 304.2M 11.4 20.3 15.9
Earth-Adapter 2.4M 49.7 50.2 50.0 (+5.9)
DINOv2 [52] (Small) Frozen 0.0M 44.3 43.3 43.8
Full Fine-Tune 22.1M 41.9 41.8 41.9
Earth-Adapter 2.4M 47.5 44.8 46.2 (+2.4)
DINOv2 [52] (Base) Frozen 0.0M 51.4 49.5 50.5
Full Fine-Tune 86.6M 46.7 49.8 49.8
Earth-Adapter 2.4M 53.0 51.8 52.4 (+1.9)
DINOv2 [52] (Large) Frozen 0.0M 57.9 49.4 53.7
Full Fine-Tune 304.2M 12.6 19.6 16.1
Earth-Adapter 2.4M 64.9 55.1 60.0 (+6.3)
Table 3: Ablation studies of Earth-Adapter on different foundation backbones. Both P2V and V2P are the DG benchmarks. Results validate that Earth-Adapter’s effectiveness on diverse backbone networks.
@@
DINOv2-L Full 304.2M 0.80 0.11 15.4
Rein 3.0M 0.68 0.12 50.0
Earth-Adapter 2.3M(1.0Msimilar-$\sim$3.9M) 0.71 0.12 59.0 (+9.0)
Figure 2: Visualization of Predicted Segmentation Maps We Compare Earth-Adapter with the Frozen DINOv2-L backbone and our baseline Rein on eight cross-domain benchmarks. For the Potsdam and Vaihingen color map, white is the Impervious surface, red is the clutter, blue is the building, color is the low vegetation, green is the tree, and yellow is the car. For LoveDA color map, red is the building, yellow is the road, blue is the water, purple is the barren, green is the forest, brown is the agriculture.
4.1 Experimental setup
Datasets and Benchmark. We conduct all experiments on several widely used remote sensing image segmentation datasets: Potsdam [49] and Vaihingen [49], both created by the International Society for Photogrammetry and Remote Sensing (ISPRS) and LoveDA [66]. The Potsdam dataset includes 38 very high resolution (VHR) True Orthophotos (TOP) images with 6,000 × 6,000 pixels. The Potsdam dataset has different imaging modes including IR-R-G channels, R-G-B channels, and R-G-B-IR channels. We use the RGB band for the following experiments. The Vaihingen dataset consists of 33 VHR TOP images with an average size of 2,494 ×2,064 pixels. The LoveDA dataset comprises a total of 4,190 images, each with a resolution of 1,024×1,024 pixels. Based on the aforementioned dataset, we follow MMSegmentation settings to split the train and val sets, crop images, and establish four benchmarks including both DA and DG: Potsdam to Vaihingen (P2V), Vaihingen to Potsdam (V2P), Rural to Urban (R2U), Urban to Rural (U2R). In our experimental setup, all images are cropped to a size of $512\times 512$.
Implementation Details. We use MMSegmentation [11] as our training and evaluation framework. We employ Mask2former [10] as our decoder, which is a highly efficient and effective algorithm for semantic segmentation tasks. For backbone, we utilize the Dinov2 [52] series models to extract features, which are subsequently utilized as input for the Mask2former. During training, we utilize AdamW [47] as the optimizer, with a learning rate of 1e-5 for the backbone, 1e-4 for the decoder, and 1e-4 for the relevant parameters in PEFT. Within the DACS training framework, we set $\alpha$ to 0.99 and $\lambda_{uda}$ to 0.5.
4.2 Cross-Domain Performance Comparison
As depicted in Table 1, we conduct a comparative analysis of the existing mainstream PEFT methods, such as Adapter [28], LoRA [31], and VPT [37], against our Earth-Adapter across eight cross-domain benchmarks,with state-of-the-art method Rein [67] in natural scene as our baseline. These benchmarks encompass four DA scenarios (P2V, V2P, R2U, and U2R) and four DG scenarios that mirror the DA ones.
Figure 4: Parameter-sensitive experiments of the cut-off frequency of DFT. The results on P2V (DA) and U2R (DG) benchmarks show Earth-Adapter’s robustness to the cut-off frequency.
Starting with the DA experiments, when compared to the performance of DA-specialized models, a significant number of existing PEFT methods exhibit performance collapse. For instance, LoRA achieves only 17.8% mIoU in the P2V (DA) benchmark when handling RS (Remote Sensing) images, a figure even lower than that of the Frozen backbone DINOv2-L. A similar trend is observed in Full Fine-Tuning, AdaptFormer, and Rein. We posit that this issue may stem from the DA self-training paradigm, which involves intensive image-mixing operations that these PEFT methods struggle to adapt to. However, two PEFT methods, Adapter and Rein, demonstrate relatively robust performance in the DA setting. When compared to these methods, our Earth-Adapter proves to be more effective in adapting to target domains. It outperforms baseline Rein by 7.5% mIoU in P2V, 3.1% mIoU in R2U, 24.0% mIoU in U2R, and 9.0% mIoU on average across the DA benchmarks. Notably, in the U2R (DA) benchmark, Earth-Adapter significantly mitigates the collapse of Rein, boosting the mIoU from 26.0% to 50.0%.
Shifting to the DG experiments, the performance collapse persists, though the trend is somewhat alleviated. In contrast to DA, the Frozen model reveals the potential generalizability of DINOv2-L. Nevertheless, FFT, LoRA, and AdaptFormer still exhibit limitations. On the other hand VPT, and Rein continue to show strong transfer capabilities, achieving 53.5% and 53.1% mIoU on average, surpassing the Frozen model by 1.7%, and 1.3% mIoU. Similar to the DA benchmark, Earth-Adapter also achieves state-of-the-art (SOTA) performance in DG, with improvements of 3.1% and 4.4% mIoU over the baseline Rein and Adapter, respectively.
4.3 Ablation Study
Analysis on Per-class Performance. In this part, we demonstrate the detailed performance of each class on four benchmarks: P2V (DA), P2V (DG), U2R (DA), and U2R (DG) between Frozen DINOv2-L, Rein, and our Earth-Adapter. All the results are shown in Table 2, and we first discuss the P2V experiment results.
In the P2V (DA), both Rein and Earth-Adapter show significant improvement on the Vegetation class compared with the Frozen DINOv2-L. Nevertheless, Earth-Adapter achieves a much higher performance than Rein, with 7.5% mIoU. A similar trend can be observed in the P2V (DG), where Earth-Adapter surpasses the other two methods by 7.0% and 4.1% mIoU respectively. Looking at the performance changes of the Vegetation and Tree classes, something interesting stands out. In the P2V (DA), both Earth-Adapter and Rein perform better on the Tree class than the Frozen. In contrast, in the P2V (DG), the performance of the Frozen with Rein on the Tree class drops drastically from 74.1% to 23.0% mIoU. As indicated in prior research [21], Rein’s features struggle to distinguish between the Vegetation and Tree classes due to their similar appearance from the RS perspective, especially in DG setting which faces more unpredictable domain shifts than DA. Moreover, in P2V (DG), comparing Rein with the Frozen, almost all other classes improve except for the Tree class. Thus, it can be empirically inferred that the improvement of the Vegetation class by Rein comes at the cost of a decline in the performance of the Tree class. However, Earth-Adapter can reduce this negative impact. It does this by optimizing the representation to tell the two classes apart and boost their performance.
Figure 5: Parameter-sensitive experiments of the projector dimension in Adapter. The results on P2V (DA) and U2R (DG) benchmarks show Earth-Adapter’s robustness to dimensions.
In the LoveDA experiments, as previously discussed, both the Frozen and Rein methods suffer from performance collapse in the U2R (DA) setting. When examining the performance at a more class level, it becomes evident that the primary decline for Rein occurs in the Barren and Agriculture classes. The performance drop in the Barren class can be attributed to it being the most sparse class within the LoveDA dataset. However, the performance of both the Frozen and Rein methods on the Agriculture class is a concerning 0.0% mIoU. In our opinion, the main factor probably is the presence of artifacts, as illustrated in Figure 1. Most PEFT methods from the CV field are unable to effectively handle the disturbances caused by these artifacts. Nevertheless, Earth-Adapter is capable of effectively utilizing the target domain data, achieving a notable 13.7% mIoU in the Barren class. Furthermore, as depicted in Figure 1, Earth-Adapter has a significant advantage with its frequency-domain divide-and-conquering ability, which enables it to improve the Agriculture performance from 0 to 63.2% mIoU. In the U2R (DG) experiments, both the Frozen and Rein methods demonstrate their effectiveness and significantly address the issues present in the DA setting. Additionally, Earth-Adapter achieves SOTA performance with improvements across most classes.
Performance on Different Backbones. Besides the DINOv2-L backbone in Table 1, Earth-Adapter is also applied to CLIP (ViT-L) and other DINOv2 scales: ViT-Small and ViT-Base, with results in Table 3. These results show Earth-Adapter improves performance across diverse backbones with varying parameter scales. Notably, Earth-Adapter achieves larger improvements on larger models, such as ViT-L, with 7.0% and 5.7% mIoU gains, more than those on small and base scale ViTs. This suggests Earth-Adapter has a greater performance upper bound.
@@ 4.3 Ablation Study
Component Effectiveness Ablation. In Table 4, ablation studies on MoA’s components and adapter numbers show that using only HF or LF adapters decreases performance compared to spatial adapters, with HF causing a larger drop. This aligns with PCA visualizations in 1 (b) and 3, where HF features have more artifacts and noise, while LF features are smooth with clear global semantics. Thus, using only HF severely damages performance. Although only using LF features outperforms HF features (50.9%mIoU), it still underperforms spatial features (56.9% mIoU) due to lacking semantic details. After that, we also conduct more experiments on the number of Spatial adapters, and reveal the best composition is the one spatial adapter with one HF and LH adapters.
Performance-Speed Trade-Off Ablation. We compare the parameters, speed, and performance between FFT, Rein, and Earth-Adapter in Table 5. Significantly, at the same time of close training and inference speed with Rein, Earth-Adapter achieves better performance (9.0% mIoU improvement), with a smaller parameter scale.
Hyper-Parameter Sensitive Ablation. In this section, we conduct an ablation study to test Earth-Adapter with two hyper-parameter variations on both the P2V (DA) and U2R (DG) benchmarks. Figure 4 presents the studies of the cut-off frequency with the dimension as 16, and Figure 5 illustrates the studies of the projector dimension in the Adapter with the cut-off frequency as 0.3. Both experiments show a consistent trend: Earth-Adapter shows greater robustness in the DA than in the DG setting. We posit that this occurs because the DA task inherently possesses higher certainty compared to DG. In the DA, even when hyper-parameters are altered, Earth-Adapter can effectively learn from the augmented target domain data, thereby maintaining performance stability. Despite DG’s uncertainty, Earth-Adapter shows a relatively stable performance in a certain parameter range. For instance, when the cut-off frequency is between 0.3 and 0.4, performance fluctuations are minimal. Similarly, for the projector dimension, performance is stable across a broad range from 16 to 256 and only fluctuates when the dimension is 512. Importantly, there are no other hyper-parameters requiring tuning. These results underscore Earth-Adapter’s superiority in handling parameter variations, offering a reliable solution for both DA and DG.
Layer Adapter Earth-Adapter
U2R (DA) R2U (DG) U2R (DA) R2U (DG)
Frozen 11.8 57.1 11.8 57.1
[0,1,2,3,4,5] 44.7 57.4 47.6 59.0
[18,19,20,21,22,23] 46.0 58.1 50.0 58.5
- - - 48.6 60.1 61.0 58.6
Full 0.3 47.0 60.9 61.3 58.1
shallow 3 0.3 48.3 60.1 63.8 58.0
shallow 6 0.3 47.6 60.1 63.8 59.0
deep 3 0.3 47.4 61.6 63.7 58.7
deep 6 0.3 48.0 60.6 64.9 58.5
deep 3 / deep 6 0.2 / 0.3 50.0 62.2 64.9 59.0
4.4 Analysis of Overall Parameter Configuration.
The detailed hyperparameter configurations for each benchmark are presented in Table 6. Overall, our method demonstrates exceptional robustness, exhibiting relatively low sensitivity to parameter variations. Below we provide an in-depth analysis of the hyperparameter effects.
@@
Effect of Frequency Adapter layer. We also conducted experiments on the configuration of frequency-domain adapters across different layers. Our study primarily focused on the Dinov2-Large model, a 24-layer Vision Transformer architecture. We evaluated four distinct adapter configurations: {0,1,2}, {0,1,2,3,4,5}, {21,22,23}, and {18,19,20,21,22,23}. As shown in Table 6, different benchmarks achieved optimal performance with varying frequency-domain adapter configuration. For instance, P2V (DA) attained its peak performance of 67.7% mIoU with adapters in 0,1,2, while P2V (DG) achieved its best results with the {18,19,20,21,22,23} configuration. Similar to the adapter dimension findings, the frequency adapter layers demonstrate task-dependent optimization patterns. Notably, three DG tasks (P2V (DG), V2P (DG), and U2R (DG)) consistently performed best with adapters in the layers {18,19,20,21,22,23}, suggesting that domain generalization (DG) requires more adapter parameters to capture general semantic representations to improve cross-domain performance.
4.5 Discussion on Adapter Layer
@@
Layer-wise comparison between Adapter and Earth-Adapter. We analyze the layer-wise behavior of Adapter and compare the performance of Adapter and Earth-Adapter under different layer configurations (Table 7). While prior work [71] suggests that deeper layers capture more semantically meaningful features, our Earth-Adapter demonstrates superior performance in both shallow-layer (first six) and deep-layer (last six) fine-tuning. For instance, in the U2R (DA) task, Earth-Adapter (shallow fine-tuning) outperforms the frozen model by 35.8% mIoU, while in R2U (DG), it surpasses the frozen baseline by 1.9% mIoU. Similar improvements are observed for deep-layer fine-tuning, achieving +38.2% mIoU (U2R (DA)) and +1.6% mIoU (U2R (DG)). Moreover, compared to Adapter, Earth-Adapter consistently shows better adaptation—+2.9% mIoU (U2R (DA), shallow) and +0.4% mIoU (R2U (DG), deep). These results highlight Earth-Adapter’s robustness and generalization capability across different layer-wise configurations.
@@
Detailed analysis of parameter configuration of Frequency Adapter. We conduct a comprehensive analysis of the impact of frequency-domain layers on model performance, as shown in Table 8. In our default setting, the cutoff frequency of Earth-Adapter is set to 0.3. Our experiments reveal that these frequency layers play a crucial role in the overall performance of Earth-Adapter. Notably, integrating the Frequency Adapter results in a slight performance drop in two benchmarks: U2R (DA) ($-1.6\%$ mIoU) and R2U (DG) ($-0.5\%$ mIoU). However, with an efficient parameter search, Earth-Adapter still surpasses the Spatial Adapter. As shown in the table, Earth-Adapter achieves 50.0% mIoU in the U2R (DA) task, outperforming Spatial-Adapter by 1.4% mIoU. Similarly, it attains 59.0% mIoU in the R2U (DG) task, exceeding Spatial Adapter by 0.4% mIoU. A key observation is that Earth-Adapter exhibits sensitivity to frequency layer configurations and cutoff frequencies under certain benchmarks—an aspect we aim to refine in future work.
5 Conclusion
In this paper, we first highlight the performance drop when transferring existing PEFT methods to the RS cross-domain experiments. Detailedly, we analyze these problems through the research [12] and our artifacts-related findings, then raise that existing methods can not effectively handle the artifacts appearing in RS images. To tackle this, we introduce Earth-Adapter, our RS-specialized PEFT method. It is a simple-yet-effective and the first PEFT method for cross-domain RS semantic segmentation. In the experiment section, we first comprehensively evaluate Earth-Adapter’s performance and compare with other methods on widely-used benchmarks. Secondly, we analyze Earth-Adapter’s effectiveness with different backbones, as well as the trade-off between parameters, speed, and performance. Finally, we visualize the decoder’s predictions and expert outputs. Our results demonstrate Earth-Adapter’s power, efficiency, robustness, and explainability.
However, there is still space for improvement. For example, it’s still unclear on which backbone layers it should be applied, and how to further enhance the utilization of frequency features. In our subsequent work, we’ll pursue more potential optimization. The code will be public soon, and we hope Earth-Adapter will be helpful to this community.

### v4

@@ Introduction
DINOv2-S
Frozen 0.0M 44.3 43.3 43.8
Earth-Adapter 2.6M~9.6M 47.5 44.8 46.2
DINOv2-B
Frozen 0.0M 51.4 49.5 50.5
Earth-Adapter 2.6M~9.6M 53.0 51.8 52.4
DINOv2-L
Frozen 0.0M 57.9 49.4 53.7
Earth-Adapter 2.6M~9.6M 64.9 55.1 60.0
RS-pretrained VFMs
Backbone Methods Params P2V(DA) V2P(DA) Avg.
MTP-L
Frozen 0.0M 32.0 33.0 32.5
Earth-Adapter 2.6M~9.6M 40.8 41.0 40.9
ScaleMAE-L
Frozen 0.0M 20.0 19.8 19.9
Earth-Adapter 2.6M~9.6M 20.2 29.6 24.9
DOFA-L
Frozen 0.0M 9.9 13.7 11.8
Earth-Adapter 2.6M~9.6M 26.2 33.2 20.7
Table 3: Ablation studies of Earth-Adapter on different VFMs. Results validate that Earth-Adapter’s effectiveness on diverse backbone networks.
Figure 3: Visualization of Predicted Segmentation Maps We Compare Earth-Adapter with the Frozen DINOv2-L backbone and our baseline Rein on eight cross-domain benchmarks. For the Potsdam and Vaihingen color map, white is the Impervious surface,red is the clutter, blue is the building, Cyan is the low vegetation, green is the tree, and yellow is the car. For LoveDA color map, red is the building, yellow is the road, blue is the water, purple is the barren, green is the forest, brown is the agriculture.
Experimental setup
Datasets and Benchmarks. We conduct all experiments on several widely used RS image segmentation datasets: Potsdam (Markus Gerke 2014) , Vaihingen (Markus Gerke 2014), LoveDA (Wang et al. 2021), and iSAID (Waqas Zamir et al. 2019). Following the standard MMSegmentation (Contributors 2020) configurations, we split each dataset into training and validation sets, apply cropping, and construct four semantic segmentation (SS) benchmarks: Potsdam(P), Vaihingen(V),LoveDA(L),iSAID(i), along with four domain adaptation (DA) and four domain generalization (DG) tasks: Potsdam to Vaihingen (P2V), Vaihingen to Potsdam (V2P), Rural to Urban (R2U), Urban to Rural (U2R). In our experimental setup, all images are cropped to a size of $512\times 512$.
Implementation Details. We use MMSegmentation as our training and evaluation framework. We employ Mask2former (Cheng et al. 2022) as our decoder, which is a highly efficient and effective algorithm for semantic segmentation tasks. For backbone, we utilize the Dinov2-Large (Oquab et al. 2023) model to extract features, which are subsequently utilized as input for the Mask2former. During training, we utilize AdamW (Loshchilov and Hutter 2019) as the optimizer, with a learning rate of 1e-5 for the backbone, 1e-4 for the decoder, and 1e-4 for the relevant parameters in PEFT. Within the DACS training framework, we set $\lambda_{da}$ to 0.5.
Main Result
As depicted in Table 1, we conduct a comparative analysis of the existing mainstream PEFT methods, such as Adapter (Houlsby et al. 2019a), LoRA (Hu et al. 2021), and VPT (Jia et al. 2022), against our Earth-Adapter across eight cross-domain benchmarks,with state-of-the-art method Rein (Wei et al. 2024) in natural scene as our baseline. These benchmarks encompass four DA scenarios (P2V, V2P, R2U, and U2R) and four DG scenarios that mirror the DA ones.
Starting with the DA experiments, many existing PEFT methods suffer from severe performance degradation when compared to DA-specialized models. For example, LoRA achieves only 17.8% mIoU on the P2V (DA) benchmark for RS images, even lower than the Frozen DINOv2-L backbone. In contrast, our Earth-Adapter demonstrates much stronger domain adaptation capabilities. It surpasses the baseline Rein by 7.5% mIoU on P2V, 3.1% mIoU on R2U, 24.0% mIoU on U2R, and achieves an average improvement of 9.0% mIoU across all DA benchmarks. Shifting to the DG experiments, the issue of performance degradation remains, although the trend is somewhat mitigated. Unlike in the DA setting, the Frozen model highlights the potential generalizability of DINOv2-L. However, FFT, LoRA, and AdaptFormer continue to show clear limitations. Consistent with the DA benchmarks, Earth-Adapter achieves state-of-the-art (SOTA) performance in DG, surpassing the baseline Rein and Adapter by 3.1% and 4.4% mIoU, respectively.
We also conducted comparisons on four SS benchmarks, with the results summarized in Table 2. Regarding PEFT-related methods, conventional approaches such as LoRA and AdaptFormer struggle to adapt to the characteristics of remote sensing images, often disrupting the representations of VFMs during fine-tuning and ultimately leading to poor performance. In contrast, our Earth-Adapter consistently achieves the best performance among all PEFT methods. We attribute this advantage to its unique design, which mitigates high-dimensional artifacts and effectively ”denoises” semantic representations, thereby enabling more accurate pixel-level feature extraction and superior semantic segmentation performance. Overall, our Earth-Adapter surpasses the baseline Rein by an average of 1.2% mIoU, highlighting its effectiveness.
Ablation and Analysis
DINOv2-L Full 304.2M 0.80 0.11 15.4
Rein 3.0M 0.68 0.12 50.0
Earth-Adapter 2.6M~9.6M 0.71 0.12 59.0 (+9.0)
@@
Ablation of Different Backbones. Beyond the DINOv2-Large backbone used in the main results (Tables 1 and 2), we further evaluated other DINOv2 variants. Results for ViT-Small and ViT-Base (Table 3, rows 3–8) show that Earth-Adapter consistently improves performance across backbones of different scales. We also tested VFMs pre-trained on remote sensing data, including MTP-Large (Wang et al. 2024), ScaleMAE-Large (Reed et al. 2023), and DOFA-Large (Xiong et al. 2024) (Table 3, rows 11–16). Earth-Adapter again delivers stable improvements in segmentation performance. Finally, we observed that DINOv2 models outperform those pre-trained on remote sensing data, likely because the latter are trained on smaller-scale datasets and thus remain less mature.
Component Effectiveness Ablation. In Table 4, ablation studies on MoA’s components and adapter numbers show that using only HF or LF adapters decreases performance compared to spatial adapters, with HF causing a larger drop. This aligns with PCA visualizations in 2 (b) and 4, where HF features have more artifacts and noise, while LF features are smooth with clear global semantics. Thus, using only HF severely damages performance. Although only using LF features outperforms HF features (50.9%mIoU), it still underperforms spatial features (56.9% mIoU) due to lacking semantic details. After that, we also conduct more experiments on the number of Spatial adapters, and reveal the best composition is the one spatial adapter with one HF and LF adapters.
@@
Method P2V(DG) V2P(DG) Avg.
w/o DR 63.6 53.2 58.4
w/ DR 64.9 55.1 60.0(+1.6)
Table 6: Ablation of Dynamic Router.
Performance-Speed Trade-Off Analysis. We compare the parameters, speed, and performance between FFT, Rein, and Earth-Adapter in Table 4. Significantly, at the same time of close training and inference speed with Rein, Earth-Adapter achieves better performance (9.0% mIoU improvement), with a smaller parameter scale. This further confirms the efficiency of Earth-Adapter, which can be attributed to its simple yet effective design methodology, making it particularly well-suited for semantic segmentation tasks on remote sensing images.
Ablation of Dynamic Router. As shown in Tab. 6, the use of the dynamic router leads to better results, yielding an average improvement of 1.6% mIoU across the two DG benchmarks. Compared with static weights (where we assign a weight of 1/3 to each static route), the dynamic router can adaptively adjust the feature allocation weights, enabling more effective representation optimization.
Conclusion
Vision Foundation Models (VFMs) excel across diverse visual tasks, and pairing them with Parameter-Efficient Fine-Tuning (PEFT) is an effective way to adapt them to downstream applications. However, existing PEFT methods often fall short on remote sensing (RS) semantic segmentation because they fail to suppress the pervasive artifacts in VFM deep features for RS imagery. To address this limitation, we introduce Earth-Adapter, a simple yet effective PEFT framework tailored for RS segmentation. It employs a frequency-guided mixture of adapters (MoA) that isolates artifacts into the high-frequency subspace, fine-tunes features within each subspace, and adaptively fuses them through a learnable router.
Acknowledgments
This work was partly supported by National Natural Science Foundation of China (62301046, 62506229) and Natural Science Foundation of Shanghai (25ZR1402268).
Appendix
Training Details
Dataset and Benchmark
Dataset Resolution Categories Train split Val Split
Potsdam $6000\times 6000$ 6 24 14
Vaihingen $\sim 2494\times 2064$ 6 16 17
LoveDA $1024\times 1024$ 7 2521 1668
iSAID $800\times 800\sim 13000\times 13000$ 16 1403 468
Table 7: Dataset overview
- - - 48.6 60.1 61.0 58.6
Full 0.3 47.0 60.9 61.3 58.1
shallow 3 0.3 48.3 60.1 63.8 58.0
shallow 6 0.3 47.6 60.1 63.8 59.0
deep 3 0.3 47.4 61.6 63.7 58.7
deep 6 0.3 48.0 60.6 64.9 58.5
deep 3 / deep 6 0.2 / 0.3 50.0 62.2 64.9 59.0
- - - 48.6 60.1 61.0 58.6
Full 0.3 47.0 60.9 61.3 58.1
shallow 3 0.3 48.3 60.1 63.8 58.0
shallow 6 0.3 47.6 60.1 63.8 59.0
deep 3 0.3 47.4 61.6 63.7 58.7
deep 6 0.3 48.0 60.6 64.9 58.5
deep 3 / deep 6 0.2 / 0.3 50.0 62.2 64.9 59.0
Lr 1e-4
Training iters 20k/40k
Batch size 4/8
$\alpha$ 0.99
$\lambda_{da}$ 0.5
Adapter dim 16/24/32/64
Cutoff frequency 0.2/0.3
Activation layer pre3/pre6/suf3/suf6
Table 10: Hyper-parameter overview
Layer Adapter Earth-Adapter
U2R (DA) R2U (DG) U2R (DA) R2U (DG)
Frozen 11.8 57.1 11.8 57.1
[0,1,2,3,4,5] 44.7 57.4 47.6 59.0
[18,19,20,21,22,23] 46.0 58.1 50.0 58.5
@@ Acknowledgments
We conduct experiments on several major optical remote sensing image segmentation datasets. The original datasets have varying resolutions, ranging from $1024\times 1024$ to $6000\times 6000$, which we uniformly crop to $512\times 512$ for training and evaluation. The detailed information of the datasets are shown in the table below.
Training Framework
In this study, we explore the application of Earth-Adapter across three distinct tasks: In-Domain(SS), DG and DA Semantic Segmentation. For SS and DG tasks, we train Earth-Adapter with Mask2former in end-to-end supervised learning as Equation 2. For DA Semantic Segmentation, we use the self-training framework introduced in DACS (Tranheden et al. 2021) without using target domain labels. The pseudo labels for target domain images are produced by a teacher network $G_{\theta,\xi,\epsilon}^{\dagger}$:
where $\cdot$] denotes the Iverson bracket, and $j\in HW$ is the spatial index of the pixel in the image. The teacher network remains inactive during the training process, with its parameters being updated based on an exponential moving average (EMA) of the student model’s parameters:
where $(\mathbf{x_{S}},\mathbf{y_{S}})$ is the data sample from the source domain, and $\mathcal{M}$ is generated by randomly selecting half of the categories based on the labels of the source domain $\mathbf{y_{S}}$. The Optimization objective for the DA task if defined by:
Hyper-Parameter Configuration
As shown in Table 7, we provide an overview of training hyperparameters. Specifically, for DA and DG tasks, we train for 20k iterations with a batch size of 4; for SS tasks, we train for 40k iterations with a batch size of 8. Notably, for the iSAID SS task, we train for 80k iterations with a batch size of 8. Other configurations are detailed in rows 5 to 7 of the table.
More Ablation and Analysis
Analysis of Overall Parameter Configuration.
The detailed hyperparameter configurations for each benchmark are presented in Table 12. Overall, our algorithm demonstrates exceptional robustness, exhibiting relatively low sensitivity to parameter variations. Below we provide an in-depth analysis of the hyperparameter effects.
@@
Effect of Frequency Adapter layer. We also conducted experiments on the configuration of frequency-domain adapters across different layers. Our study primarily focused on the Dinov2-Large model, a 24-layer Vision Transformer architecture. We evaluated four distinct adapter configurations: [0,1,2], [0,1,2,3,4,5], [21,22,23], and [18,19,20,21,22,23]. As shown in Table 12, different benchmarks achieved optimal performance with varying frequency-domain adapter configuration. For instance, P2V (DA) attained its peak performance of 67.7% mIoU with adapters in [0,1,2], while P2V (DG) achieved its best results with the [18,19,20,21,22,23] configuration. Similar to the adapter dimension findings, the frequency adapter layers demonstrate task-dependent optimization patterns. Notably, three DG tasks (P2V (DG), V2P (DG), and U2R (DG)) consistently performed best with adapters in the layers [18,19,20,21,22,23], suggesting that domain generalization (DG) requires more adapter parameters to capture general semantic representations to improve cross-domain performance.
@@
Layer-wise comparison between Adapter and Earth-Adapter. We analyze the layer-wise behavior of Adapter and compare the performance of Adapter and Earth-Adapter under different layer configurations (Table 11). While prior work (Oquab et al. 2023) suggests that deeper layers capture more semantically meaningful features, our Earth-Adapter demonstrates superior performance in both shallow-layer (first six) and deep-layer (last six) fine-tuning. For instance, in the U2R (DA) task, Earth-Adapter (shallow fine-tuning) outperforms the frozen model by 35.8% mIoU, while in R2U (DG), it surpasses the frozen baseline by 1.9% mIoU. Similar improvements are observed for deep-layer fine-tuning, achieving +38.2% mIoU (U2R (DA)) and +1.6% mIoU (U2R (DG)). Moreover, compared to Adapter, Earth-Adapter consistently shows better adaptation—+2.9% mIoU (U2R (DA), shallow) and +0.4% mIoU (R2U (DG), deep). These results highlight Earth-Adapter’s robustness and generalization capability across different layer-wise configurations.
@@
Detailed analysis of parameter configuration of Frequency Adapter. We conduct a comprehensive analysis of the impact of frequency-domain layers on model performance, as shown in Table 9. In our default setting, the cutoff frequency of Earth-Adapter is set to 0.3. Our experiments reveal that these frequency layers play a crucial role in the overall performance of Earth-Adapter. Notably, integrating the Frequency Adapter results in a slight performance drop in two benchmarks: U2R (DA) ($1.6\%$ mIoU) and R2U (DG) ($0.5\%$ mIoU). However, with an efficient parameter search, Earth-Adapter still surpasses the Spatial-Adapter. As shown in the table, Earth-Adapter achieves 50.0% mIoU in the U2R (DA) task, outperforming Spatial-Adapter by 1.4% mIoU. Similarly, it attains 59.0% mIoU in the R2U (DG) task, exceeding Spatial-Adapter by 0.4% mIoU. A key observation is that Earth-Adapter exhibits sensitivity to frequency layer configurations and cutoff frequencies under certain benchmarks—which we aim to refine in future work.